# Learnable Infinite Taylor Gaussian for Dynamic View Rendering

### v1

@@ Appendix B Performance Analysis of Large-Scale Data Scene Reconstruction
To comprehensively evaluate the performance of our model, we conducted tests using the Technicolor Light Field Dataset, a large-scale and high-quality visual resource. This dataset captures video sequences through a 4X4 array of 16 cameras that are precisely synchronized in time, enabling the acquisition of high spatial resolution images with a resolution of up to 2048X1088 pixels. The Technicolor Light Field Dataset encompasses five unique scenes: Birthday, Painter, Theater, Train, and Fabien. These scenes not only enhance the dataset’s diversity but also provide a broad testing environment for model evaluation. We selected two scenes, Birthday and Painter, for quantitative and qualitative analysis.
Our evaluation approach integrates both quantitative and qualitative dimensions. For quantitative analysis, we employed a series of standardized metrics, including PSNR, SSIM, and LPIPS, to objectively measure the model’s performance across different scenarios. These metrics offer a clear perspective on the model’s capabilities in various testing conditions. For qualitative analysis, we delved deeper into the model’s visual performance when handling complex scenes. This included assessing the model’s ability to capture details and reconstruct object surface textures. By visually comparing the model’s outputs with real-world scenes, we gained a deeper understanding of the model’s potential and limitations in practical applications.
As shown in Table 4, our algorithm outperforms other methods, delivering superior results across multiple standardized metrics. Specifically, our model excels in PSNR, SSIM, and LPIPS, demonstrating its effectiveness in accurately evaluating performance across various scenarios. To vividly illustrate the differences between our method and others, we visualized the qualitative results. As highlighted in the red-boxed regions, it is evident that our method can render higher-quality images. As shown in Figure 67, we can see that in the Birthday scene, our reconstruction captures details better compared to other models. Several issues are observed in the reconstructions of 4DGS and FSGS: in the area marked by the blue box, both methods exhibit reconstruction blurriness; in the area marked by the red box, neither 4DGS nor FSGS successfully reconstructs the yellow object near the person’s nose bridge, and the images generated by both methods have relatively lower resolution. Additionally, FSGS introduces motion blur artifacts. In the area marked by the yellow box, both methods make errors in reconstruction, mistakenly generating a red object beneath the green leaf Lastly, in the area marked by the green box, both 4DGS and FSGS fail to accurately reconstruct the text along the edges.

### v2

@@ Appendix B Performance Analysis of Large-Scale Data Scene Reconstruction
To better evaluate the performance of Large-Scale Data Scene Reconstruction, we conducted a qualitative analysis on the Technicolor Light Field Dataset. For a more in-depth assessment, we explored the model’s visual performance in handling complex scenes, focusing on its ability to capture fine details and reconstruct object surface textures. By visually comparing the model’s outputs with real-world scenes, we gained deeper insights into its strengths and limitations in practical applications.
As highlighted in the boxed regions, it is evident that our method can render higher-quality images. As shown in Figure 56, we can see that in the Birthday scene, our reconstruction captures details better compared to other models. Several issues are observed in the reconstructions of 4DGS and FSGS: in the area marked by the blue box, both methods exhibit reconstruction blurriness; in the area marked by the red box, neither 4DGS nor FSGS successfully reconstructs the yellow object near the person’s nose bridge, and the images generated by both methods have relatively lower resolution. Additionally, FSGS introduces motion blur artifacts. In the area marked by the yellow box, both methods make errors in reconstruction, mistakenly generating a red object beneath the green leaf Lastly, in the area marked by the green box, both 4DGS and FSGS fail to accurately reconstruct the text along the edges.